# Chaos Engineering: A Multi-Vocal Literature Review

### v1

@@ 4.2. Socio-Technical Challenges
Chaos engineering can address various socio-technical challenges, expanding its impact beyond technical issues to encompass the complex interplay between people, processes, and technology (Miles, 2019b; Thomas et al., 2019; Patriarca et al., 2018).
@@
Effective risk management is critical in both production and non-production environments, yet many systems are exposed to unexpected failures due to insufficient preparation (Bremmers, 2021; Naqvi et al., 2022). In non-production settings, such as development or staging, failure scenarios are often inadequately tested, leading to a false sense of security (Wickramasinghe, 2023; Jernberg et al., 2020). These environments typically lack the complexity of live production systems, meaning that when software is deployed, it may not be fully equipped to handle adverse real-world events (Green, 2023). Consequently, when systems reach production, where they face real user traffic, external dependencies, and unpredictable conditions, even minor issues can escalate to significant outages, resulting in costly downtime and degraded user experiences (Gill, 2021; Zhang et al., 2019; Ji et al., 2023). Chaos engineering solves these challenges by enabling controlled experimentation that tests system behavior under simulated and real-world failure conditions (Torkura et al., 2020; Kostic, 2024). In production environments, controlled blast radiuses ensure that failures are confined to isolated parts of the system, thereby preventing widespread disruption (Palani and Gupta, 2023; Siwach et al., 2022; Gianchandani et al., 2022; Mooney, 2023; Bocetta, 2019). Predefined abort conditions and rollback procedures allow teams to stop experiments and recover quickly if the stability of the system is threatened (Abdul, 2024; Sbiai et al., 2023; Siwach et al., 2022).
@@
In complex, distributed systems, incident management faces significant challenges due to the unpredictability of failures and the limitations of reactive strategies (Gianchandani et al., 2022). In these approaches, teams respond to incidents only after they occur, focusing on diagnosing and resolving issues after disruptions have already impacted operations (Bremmers, 2021). This often leads to prolonged downtime and repeated failures, as reactive methods limit the opportunities to address these problems proactively (Patel, 2022). Chaos engineering offers a proactive approach by simulating targeted failures that stress critical parts of the system, revealing weaknesses in the incident response process (Anwar and Arunachalam, 2019; Siwach et al., 2022). For example, simulating a sudden network partition might show whether monitoring systems detect the issue promptly and if the system can reroute traffic as intended (Vinisky, 2024; Green, 2023; Gill, 2021). Similarly, inducing a Central Processing Unit (CPU) or memory exhaustion event could expose delays in triggering alerts or inefficiencies in the automated recovery process (Camacho et al., 2022; Kadikar, 2023; Palani and Gupta, 2023; Roa, 2022). By creating these controlled failure scenarios, chaos engineering helps teams identify where their incident response protocols fall short, whether in monitoring, alert escalation, or failover processes (Wu, 2021; Zhang et al., 2021a). This enables teams to proactively adjust their strategies, improving their ability to handle actual incidents faster, more accurately, and with minimal disruption (Dua, 2024; Rivera et al., 2023).
@@ 4.3. Benefits
Table 9 outlines the organizational benefits derived from the functionalities described in Section 3.2, addressing the previously identified challenges. These functions include defining chaos experiments, facilitating team communication, notifying stakeholders, maintaining documentation, and using real-time monitoring dashboards. The corresponding activities, such as establishing experimental parameters, analyzing system performance metrics, and fostering cross-functional collaboration, contribute to several organizational advantages. These include improved system understanding, enhanced error-handling capabilities, vital team coordination during stress conditions, and proactive system tuning. Furthermore, automation of tasks and real-time system monitoring can minimize disruptions and optimize system performance and user experience.
Table 9. Mapping of Key Activities Performed by Chaos Engineering with Organizational Benefits
Key Functions Activity Performed Benefits of Organization Articles Total
Define Chaos Experiments Set experiment scope, define normal conditions, hypothesize, and inject fault. Gain system insights, refine the design, establish a baseline, and test error handling. (Kostic, 2024; Gill, 2021; Manna, 2021; Inc., 2023; Wickramasinghe, 2023; Mooney, 2023; Fawcett, 2020; Vinisky, 2024; Al-Said Ahmad et al., 2024; Pierce et al., 2021; Torkura et al., 2020; Vu et al., 2022; Bedoya et al., 2023; Siwach et al., 2022; Zhang et al., 2019; Le, 2022; Palani and Gupta, 2023; Soldani and Brogi, 2021; Azure Readiness, 2023; Dedousis et al., 2023; Colyer, 2019; Patel, 2022; Shah, 2021; Green, 2023; Fogli et al., 2023; National Australia Bank, 2020; Cotroneo et al., 2022; Konstantinou et al., 2021; Zhang et al., 2023; Bairyev, 2023) 30
Facilitate Team Coordination Create groups for team collaboration. Strengthens coordination and insight sharing. (Gill, 2021; Manna, 2021; Dua, 2024; Kostic, 2024; Katirtzis, 2022; Krishnamurthy, 2021; Torkura et al., 2020; Green, 2023; Frank et al., 2021b; Kesim et al., 2020; Bocetta, 2019; Bremmers, 2021) 12
Notify Stakeholders Update stakeholders on progress and results. Promote transparency and prompt updates. (Gill, 2021; Manna, 2021; Dua, 2024; Kostic, 2024; Katirtzis, 2022; Krishnamurthy, 2021; Torkura et al., 2020; Green, 2023; Frank et al., 2021b; Kesim et al., 2020; Bocetta, 2019; Bremmers, 2021) 12
Maintain Knowledge Repositories Share findings for organizational learning. Create a knowledge base to analyze system stress behavior. (Gill, 2021; Manna, 2021; Dua, 2024; Kostic, 2024; Katirtzis, 2022; Krishnamurthy, 2021; Torkura et al., 2020; Green, 2023; Frank et al., 2021b; Kesim et al., 2020; Bocetta, 2019; Starr et al., 2022; Vinisky, 2024; Anwar and Arunachalam, 2019; Kadikar, 2023; Patel, 2022; Vu et al., 2022; Zhang et al., 2019, 2023; Rivera et al., 2023; Bremmers, 2021) 21
Provide Real-Time Dashboards for Visualization Track faults and visualize performance under stress. Assess system health for rapid anomaly detection and refinement. (Tivey et al., 2019; Kadikar, 2023; Pierce et al., 2021; Zhang et al., 2021a; Wickramasinghe, 2023; Gill, 2022; National Australia Bank, 2020; Simonsson et al., 2021; Al-Said Ahmad et al., 2024; Torkura et al., 2020; Gill, 2021; Kostic, 2024; Fogli et al., 2023; Bedoya et al., 2023; Manna, 2021; Krivas and Portela, 2020; Poltronieri et al., 2022; Camacho et al., 2022; Torkura et al., 2021; Pulcinelli et al., 2023; Chen et al., 2022; Yu et al., 2021; Gianchandani et al., 2022; Palani and Gupta, 2023; Anwar and Arunachalam, 2019; Krishnamurthy, 2021; Roa, 2022; Vinisky, 2024; Siwach et al., 2022; Green, 2023; Hawkins, 2020; Vu et al., 2022; Ma et al., 2023; Rivera et al., 2023) 34
Analyze Impact and Identify Root Causes of Faults Assess outcomes to measure impact and analyze root causes. Enhance resilience, identify weaknesses, and improve stability. (Bairyev, 2023; Siwach et al., 2022; Fogli et al., 2023; Zhang et al., 2019; Inc., 2023; Chen et al., 2024; Bhaskar, 2022; Al-Said Ahmad et al., 2024; Gill, 2022; Wickramasinghe, 2023; Yu et al., 2021; Ji et al., 2023; Gianchandani et al., 2022; Nikolaidis et al., 2021; Rivera et al., 2023; Konstantinou et al., 2021; Torkura et al., 2020) 17
Implement Safety Controls and Automated Rollbacks Set up safety mechanisms, automated rollbacks, and abort thresholds to prevent severe disruptions. Minimizes risk, ensures safe experimentation, and reduces downtime with quick recovery measures. (Wickramasinghe, 2023; Siwach et al., 2022; Bocetta, 2019; Gill, 2022; Fawcett, 2020; Manna, 2021; Gianchandani et al., 2022; Bairyev, 2023; Malik et al., 2023; Inc., 2023; Ikeuchi et al., 2020; Ma et al., 2023; Nikolaidis et al., 2023; Bhaskar, 2022; Simonsson et al., 2021; Gill, 2021; Krishnamurthy, 2021; Torkura et al., 2020; Katirtzis, 2022; Bedoya et al., 2023; Kadikar, 2023; Patel, 2022; Soldani and Brogi, 2021; Frank et al., 2021b; Rivera et al., 2023; Chen et al., 2024; Zhang et al., 2021a, 2019; Chen et al., 2022; Basiri et al., 2019; Bremmers, 2021; Jernberg et al., 2020; Kostic, 2024) 31
Ensure Regulatory Compliance Align chaos experiments with regulatory standards and monitor compliance. Reduce non-compliance risk and build trust by adhering to industry standards. (Wickramasinghe, 2023; Siwach et al., 2022; Bocetta, 2019; Gill, 2022; Fawcett, 2020; Manna, 2021; Gianchandani et al., 2022; Bairyev, 2023; Malik et al., 2023; Inc., 2023; Ikeuchi et al., 2020; Ma et al., 2023; Nikolaidis et al., 2023; Bhaskar, 2022; Simonsson et al., 2021; Gill, 2021; Jernberg et al., 2020; Kadikar, 2023; Kesim et al., 2020; Zhang et al., 2019; Krivas and Portela, 2020; Sbiai et al., 2023; Mooney, 2023; Dua, 2024; Torkura et al., 2021; Palani and Gupta, 2023; Pulcinelli et al., 2023; Kostic, 2024) 28
Integrate with CI/CD and Third-Party Tools Embed chaos experiments in CI/CD with monitoring integration. Automate resilience testing in the software lifecycle to boost efficiency. (Manna, 2021; Azure Readiness, 2023; Konstantinou et al., 2021; Palani and Gupta, 2023; Kostic, 2024; Torkura et al., 2021) 6
Automate Repetitive and Manual Tasks Automate setup and cleanup of chaos experiments using scripts and tools. Reduce time and errors by automating repetitive tasks for consistency and efficiency. (Camacho et al., 2022; Kadikar, 2023; Miles, 2019a; Green, 2023; Nikolaidis et al., 2021; Poltronieri et al., 2022; Fogli et al., 2023; Torkura et al., 2020; Gill, 2021; Cotroneo et al., 2022; Patel, 2022; Bairyev, 2023; Colyer, 2019; Manna, 2021; Zhang et al., 2019; Kostic, 2024; Fawcett, 2020; Jakkaraju, 2020; Zhang et al., 2021a; Al-Said Ahmad et al., 2024; Mooney, 2023; Kesim et al., 2020; Zhang et al., 2023; Jernberg et al., 2020; Bremmers, 2021; Torkura et al., 2021; Vinisky, 2024; Chen et al., 2022; Vu et al., 2022) 29
@@ 5. Chaos Engineering Framework: Taxonomy, Tools, Practices, and Evaluation (RQ3)
This section presents a taxonomy for chaos engineering and compares existing tools using the taxonomy dimensions. We also present the best practices for adopting chaos engineering in organizations.
5.1. Taxonomy of Chaos Engineering
Figure 9 shows the proposed taxonomy, which organizes the capabilities of a chaos engineering platform across several key dimensions: execution environment, automation strategy, automation mode, deployment type, and evaluation approach. This taxonomy provides organizations with a structured approach to adopting chaos engineering practices by offering clear guidelines for selecting appropriate execution environments, choosing automation strategies that integrate essential tools such as monitoring, logging, and observability, and determining the optimal deployment types (e.g., production or pre-production) based on their risk tolerance. It also helps define task modes (manual, semi-automated, or fully automated). It guides the selection of fault injection tools, ensuring that organizations have the necessary evaluation methods to measure the impact of experiments and derive actionable insights for system improvement.

### v2

@@ 5.2. Socio-Technical Challenges
Chaos engineering can address various socio-technical challenges, expanding its impact beyond technical issues to encompass the complex interplay between people, processes, and technology (Miles, 2019b; Thomas et al., 2019; Patriarca et al., 2018). For instance, the theme ‘Inconsistent Experiment Design’ maps to codes such as ‘variable test timing’ and ‘lack of standardization’. Additional examples and detailed code-theme mappings can be found in the online appendix(see Section 3.6).
@@
Effective risk management is critical in both production and non-production environments, yet many systems are exposed to unexpected failures due to insufficient preparation (Bremmers, 2021; Naqvi et al., 2022). In non-production settings, such as development or staging, failure scenarios are often inadequately tested, leading to a false sense of security (Wickramasinghe, 2023; Jernberg et al., 2020). These environments typically lack the complexity of live production systems, meaning that when software is deployed, it may not be fully equipped to handle adverse real-world events (Green, 2023). Consequently, when systems reach production, where they face real user traffic, external dependencies, and unpredictable conditions, even minor issues can escalate to significant outages, resulting in costly downtime and degraded user experiences (Gill, 2021; Zhang et al., 2019; Ji et al., 2023). Chaos engineering solves these challenges by enabling controlled experimentation that tests system behavior under simulated and real-world failure conditions (Torkura et al., 2020; Kostic, 2024). In production environments, controlled blast radiuses ensure that failures are confined to isolated parts of the system, thereby preventing widespread disruption (Palani and Gupta, 2023a; Siwach et al., 2022; Gianchandani et al., 2022; Mooney, 2023; Bocetta, 2019). Predefined abort conditions and rollback procedures allow teams to stop experiments and recover quickly if the stability of the system is threatened (Abdul, 2024; Sbiai et al., 2023; Siwach et al., 2022).
@@
In complex, distributed systems, incident management faces significant challenges due to the unpredictability of failures and the limitations of reactive strategies (Gianchandani et al., 2022). In these approaches, teams respond to incidents only after they occur, focusing on diagnosing and resolving issues after disruptions have already impacted operations (Bremmers, 2021). This often leads to prolonged downtime and repeated failures, as reactive methods limit the opportunities to address these problems proactively (Patel, 2022). Chaos engineering offers a proactive approach by simulating targeted failures that stress critical parts of the system, revealing weaknesses in the incident response process (Anwar and Arunachalam, 2019; Siwach et al., 2022). For example, simulating a sudden network partition might show whether monitoring systems detect the issue promptly and if the system can reroute traffic as intended (Vinisky, 2024; Green, 2023; Gill, 2021). Similarly, inducing a Central Processing Unit (CPU) or memory exhaustion event could expose delays in triggering alerts or inefficiencies in the automated recovery process (Camacho et al., 2022; Kadikar, 2023; Palani and Gupta, 2023a; Roa, 2022). By creating these controlled failure scenarios, chaos engineering helps teams identify where their incident response protocols fall short, whether in monitoring, alert escalation, or failover processes (Wu, 2021; Zhang et al., 2021a). This enables teams to proactively adjust their strategies, improving their ability to handle actual incidents faster, more accurately, and with minimal disruption (Dua, 2024; Rivera et al., 2023).
Table 11. Mapping of Core Chaos Engineering Functionalities with Organizational Benefits.
Core Functionality Activity Performed Benefits to Organization Articles Referenced Total
Identify Test Targets and Failure Scenarios Select critical components (e.g., APIs, services) and likely fault types based on architecture and risk. Focuses efforts on the most vulnerable areas, improving preparedness and reducing downtime (Bremmers, 2021; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Inc., 2023; Gill, 2021; Wickramasinghe, 2023; Zhang et al., 2021b; Cotroneo et al., 2022; Poltronieri et al., 2022; Zhang et al., 2021a; Frank et al., 2021a; Pierce et al., 2021; Soldani and Brogi, 2021; Al-Said Ahmad et al., 2024; Vu et al., 2022; Nikolaidis et al., 2023; Moon, 2022; Anwar and Arunachalam, 2019; Dua, 2024; Gill, 2022; Fawcett, 2020; Manna, 2021; Gianchandani et al., 2022; Abdul, 2024; Azure Readiness, 2023; Roa, 2022; Krivas and Portela, 2020; Kostic, 2024; Vinisky, 2024; Colyer, 2019; Green, 2023) 33
Define Steady-State Behavior Establish baseline performance metrics (e.g., latency, throughput) Provides a reference for detecting anomalies and validating system health under failure. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Cotroneo et al., 2022; Nikolaidis et al., 2021; Pierce et al., 2021; Frank et al., 2023; Konstantinou et al., 2021; Naqvi et al., 2022; Kesim et al., 2020; Park et al., 2019; Zhang et al., 2023; Siwach et al., 2022; Jernberg et al., 2020; Jakkaraju, 2020; Le, 2022; Moon, 2022; Bocetta, 2019; Palani and Gupta, 2023b; Gill, 2022; Fawcett, 2020; Bairyev, 2023; Abdul, 2024; Roa, 2022; Aharon, 2024; Krivas and Portela, 2020; Kostic, 2024) 33
Formulate Testable Hypotheses Describe expected system responses to faults in measurable terms. Guides experiment design and enables structured interpretation of results (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Cotroneo et al., 2022; Frank et al., 2023; Konstantinou et al., 2021; Naqvi et al., 2022; Kesim et al., 2020; Park et al., 2019; Zhang et al., 2023; Siwach et al., 2022; Jernberg et al., 2020; Jakkaraju, 2020; Bocetta, 2019; Palani and Gupta, 2023b; Gill, 2022; Fawcett, 2020; Bairyev, 2023; Abdul, 2024) 25
Inject Faults Based on Defined Scenarios Execute fault (e.g., CPU overload, network delay) on selected systems. Validates resilience strategies; exposes weak points under controlled stress (Zhang et al., 2019, 2021b; Cotroneo et al., 2022; Zhang et al., 2021a; Nikolaidis et al., 2021; Frank et al., 2021b; Pierce et al., 2021; Soldani and Brogi, 2021; Vu et al., 2022; Chen et al., 2024; Nikolaidis et al., 2023; Ma et al., 2023; Meiklejohn et al., 2021; Sondhi et al., 2021; Ikeuchi et al., 2020; Chen et al., 2022; Basiri et al., 2019; Gill, 2022; Gianchandani et al., 2022; Sbiai et al., 2023; Kadikar, 2023; Bhaskar, 2022) 22
Apply Safety Mechanisms Enforce abort conditions, rollback procedures, and limit blast radius. Minimizes impact, safeguards critical systems, ensures quick recovery. (Bremmers, 2021; Patel, 2022; Gill, 2021; Wickramasinghe, 2023; Frank et al., 2021b; Pierce et al., 2021; Soldani and Brogi, 2021; Al-Said Ahmad et al., 2024; Vu et al., 2022; Konstantinou et al., 2021; Ikeuchi et al., 2020; Chen et al., 2022; Naqvi et al., 2022; Jakkaraju, 2020; Starr et al., 2022; Katirtzis, 2022; Dua, 2024; Tivey et al., 2019; Palani and Gupta, 2023b; Bairyev, 2023; Sbiai et al., 2023; Kadikar, 2023; Green, 2023; Vinisky, 2024) 24
Notify Stakeholders During Experiment Lifecycle Send alerts and updates before, during, and after chaos experiments. Promotes transparency, readiness, and coordinated response among teams. (Torkura et al., 2020; Zhang et al., 2019; Simonsson et al., 2021; Patel, 2022; Gill, 2021; Wickramasinghe, 2023; Zhang et al., 2021b; Torkura et al., 2021; Nikolaidis et al., 2021; Pierce et al., 2021; Soldani and Brogi, 2021; Konstantinou et al., 2021; Siwach et al., 2022; Krishnamurthy, 2021; Durai, 2022; Tomka, 2024; Dua, 2024; Tivey et al., 2019; Manna, 2021; Abdul, 2024; Green, 2023; Kostic, 2024; Vinisky, 2024; Azure Readiness, 2023) 24
Capture Live System Telemetry Monitor and record real-time metrics, logs, and traces during injection. Enables detailed analysis of system and supports rapid anomaly detection. (Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Inc., 2023; Gill, 2021; Camacho et al., 2022; Poltronieri et al., 2022; Nikolaidis et al., 2021; Vu et al., 2022; Chen et al., 2024; Nikolaidis et al., 2023; Ma et al., 2023; Sondhi et al., 2021; Pulcinelli et al., 2023; Naqvi et al., 2022; Zhang et al., 2023; Siwach et al., 2022; Jernberg et al., 2020; Le, 2022; Butow, 2018) 20
Visualize Experiment Outcomes and Behavior Display telemetry on dashboards to highlight key deviations and trends. Provides insights for decisions and facilitates shared understanding. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Inc., 2023; Bremmers, 2021; Zhang et al., 2021b; Cotroneo et al., 2022; Nikolaidis et al., 2021; Green, 2023) 11
Analyze Impact & Root Causes Compare metrics to baselines and trace root causes. Identifies weaknesses, guides improvements. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Inc., 2023; Bremmers, 2021; Zhang et al., 2021b; Cotroneo et al., 2022; Poltronieri et al., 2022; Zhang et al., 2021a; Nikolaidis et al., 2021; Pierce et al., 2021; Soldani and Brogi, 2021; Bairyev, 2023; Green, 2023) 16
Archive and Document Learnings Store findings, root causes, and lessons learned in accessible knowledge bases. Preserves knowledge and supports future analysis, audits, and training. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Al-Said Ahmad et al., 2024; Vu et al., 2022; Frank et al., 2023; Chen et al., 2022; Kesim et al., 2020; Basiri et al., 2019; Camilli et al., 2022; Yu et al., 2021; Siwach et al., 2022; Konstantinou et al., 2021; Wu, 2021; Kadikar, 2023; Green, 2023) 22
Automate Chaos Testing Pipelines Schedule and execute experiments with scripts or workflows. Enhances consistency, lowers overhead, supports continuous validation. (Torkura et al., 2020; Zhang et al., 2019; Simonsson et al., 2021; Gill, 2021; Nikolaidis et al., 2021; Pierce et al., 2021; Ikeuchi et al., 2020; Chen et al., 2022; Bedoya et al., 2023; Park et al., 2019; Siwach et al., 2022; Manna, 2021; Green, 2023) 13
Integrate with Platform Tools Link chaos tools to CI/CD, monitoring, and cloud systems for orchestration. Embeds resilience testing into delivery pipelines and ensures synchronized operations. (Torkura et al., 2020; Dedousis et al., 2023; Bremmers, 2021; Cotroneo et al., 2022; Camacho et al., 2022; Poltronieri et al., 2022; Nikolaidis et al., 2021; Frank et al., 2021b; Soldani and Brogi, 2021) 9
@@ 5.3. Benefits
Table 11 outlines the organizational benefits derived from the functionalities described in Section 4.2, addressing the previously identified challenges. These functionalities were identified through our thematic analysis process. They include defining chaos experiments, facilitating team communication, notifying stakeholders, maintaining documentation, and using real-time monitoring dashboards. The corresponding activities, such as establishing experimental parameters, analyzing system performance metrics, and fostering cross-functional collaboration, contribute to several organizational advantages. These include improved system understanding, enhanced error-handling capabilities, vital team coordination during stress conditions, and proactive system tuning. Furthermore, automating tasks and monitoring systems in real-time can minimize disruptions and optimize system performance and user experience.
@@ 6. Chaos Engineering Framework: Taxonomy, Tools, Practices, and Evaluation (RQ3)
This section addresses (RQ3) by outlining proposed solutions in chaos engineering. It presents a taxonomy of tools and techniques (RQ3.1), highlights best and bad adoption practices (RQ3.2), and describes the methods used to evaluate chaos engineering approaches (RQ3.3).
6.1. Taxonomy of Chaos Engineering (RQ3.1)
To address RQ3.1, we propose a taxonomy (Figure 8) that organizes chaos engineering platforms across several key dimensions: execution environment, automation strategy, automation mode, deployment type, and evaluation approach. The dimensions of this taxonomy were developed from the core themes identified through our coding and thematic analysis. Table 12 presents an excerpt of the intermediate coding structure that grounds this taxonomy, mapping each conceptual category to its subcodes and supporting sources. This taxonomy provides organizations with a structured approach to adopting chaos engineering practices by offering clear guidelines for selecting appropriate execution environments, choosing automation strategies that integrate essential tools such as monitoring, logging, and observability, and determining the optimal deployment types (e.g., production or pre-production) based on their risk tolerance. It also helps define task modes (manual, semi-automated, or fully automated). Additionally, it guides the selection of chaos engineering tools by aligning them with specific combinations of these dimensions—for example, tools differ in aspects such as the types of faults they inject and the level of automation they support, ensuring that organizations have the necessary evaluation methods to measure the impact of experiments and derive actionable insights for system improvement.
Figure 8. A Taxonomy of Chaos Engineering Platforms and Their Capabilities.
[…]
A taxonomy of chaos engineering platforms.
Table 12. Representative Codes and Subcodes for the Taxonomy of Chaos Engineering Platforms and Their Capabilities.
Code Sub-code Sub-code
Execution Environment Virtual Machines Variability introduced by different guest OS configurations.
Resource contention issues examined under CPU/memory saturation scenarios.
Automation Mode Manual Tends to be time and resource intensive, as each step must be performed and monitored manually.
Experiments are custom designed for specific systems or hypotheses.
Carries risk of human error, especially in timing, targeting, or recovery.
Automation Mode Semi-Automated Engineer-Defined Parameters.
Human-in-the-Loop Oversight.
Automation Strategy Custom Code Used to simulate specific failure types (e.g., service crash, DB outage).
Custom scripts extend existing chaos tools to support advanced scenarios.
Allows precise control over failure timing, scope, and recovery validation.
Automation Strategy Prioritization Critical Component Identification.
Resource Allocation Strategy.
Automation Strategy Workflow Automation Ensures repeatability and reduces manual error.
Supports multi-step or multi-system coordination.
Deployment Type Production Live Environment Testing.
Real-User Interaction Observation.
Deployment Type Pre-production Complement to Production Testing.
Staging Environment Simulation.